Find the differential in y=3√6x



1
Expert's answer
2021-02-24T07:58:24-0500

"\bold {Answer}"

"y^{'} = \dfrac {3\sqrt{6}}{2\sqrt{x}} = \dfrac {3\sqrt{6x}}{2x}"


"\bold {Solution }"

"y = 3\sqrt{6x}"


"= 3\sqrt{6} × \sqrt{x}"


"y^{'} = \dfrac {d}{dx}(3\sqrt{6x})"

"= 3\sqrt{6}\dfrac {d}{dx}(\sqrt{x})"


"= 3\sqrt{6}\dfrac {d}{dx}(x^{\small\dfrac {1}{2}})"


"= \dfrac {1}{2} × 3\sqrt{6} × x ^{\small-\dfrac{1}{2}})"


"= \dfrac {3\sqrt{6}}{2\sqrt{x}}" Or "\dfrac {3\sqrt {6x}}{2x}"